OpenStudy (anonymous):

trey is taking out a loan for 85000. It is a 20 year load with an apr of 5.85% what will his monthly payment be?

OpenStudy (mathmate):

Use A=P(1+r/12)^n*(r/12)/((1+r/12)^n-1) A=monthly amount P=principal r=apr (0.0585) n=no. of periods (20*12=240) I get 601.63 a month It seems to me that you have numerous problems of this kind. It would help if you could write down the above formula and apply for your next problems.

OpenStudy (anonymous):

THANKS I DIDNT KNOW WHAT FORMULA TO USE THANKS SO MUCH

OpenStudy (mathmate):

You're welcome!